OpenStudy (superhelp101):

Solve for x. 7^(x-2)=12^(x)

OpenStudy (superhelp101):

@e.mccormick @mathmale @jim_thompson5910 @saifoo.khan can any of you guys help me ? :(

OpenStudy (superhelp101):

@GIR__

OpenStudy (superhelp101):

@shamil98

OpenStudy (superhelp101):

@Compassionate @anonymous_user @mathmale plzzzz :))

OpenStudy (superhelp101):

anyone ?

OpenStudy (anonymous_user):

I found this: 7^(x-2)=12^(x) ln(7^(x-2))=ln(12^(x)) xln(7)-2ln(7)=ln(12^(x)) xln(7)-2ln(7)-xln(12)=0 x(ln(7)-ln(12))=2ln(7) x=(2ln(7))/(ln(7)-ln(12)) x=3.8918

OpenStudy (superhelp101):

yes but I plug it back in and it doesn't work :/

OpenStudy (superhelp101):

if you don't know, no worries :)

OpenStudy (anonymous_user):

^Did that link help?

OpenStudy (superhelp101):

a little bit..I plugged it back in it works. But I was looking for an explanation like the one you sent me :) it's fine :)

OpenStudy (anonymous_user):

Sorry I couldn't explain